Wednesday, February 15, 2017

From HallCWiki
Jump to: navigation, search

Meeting Information

  • Location CC L102, 11 am
  • DocDB

Blue Jeans Connection Information

  1. Enter Meeting ID: 654704307
  2. Press #

Agenda

  1. Review systems owners and update information if needed
  2. Point system owners to Wiki's that need to be addressed for the upcoming commissioning run
  3. Address system specific points of interest for the upcoming commissioning run

System Owners

HMS

SHMS

Minutes

General

  • All groups are requested to update documentation on the wiki link
    • We will use these as references for shift crew later.
    • See links under System Owners
    • Use the HMS Drift Chambers entries as an example
    • Links should point to documents in the DocDB
    • Don't just upload to the wiki
  • Please ensure any hardware changes you make get put in the HCLog
    • Even if they shouldn't matter to others, it is very useful to have the history logged *AND* a record of where people have been working
      • Sometimes another group's cable gets knocked loose, etc.

Miscellaneous

  • Gunning: Raster pickoffs available
    • Need to get them in to the DAQ
  • S. Wood: V2F hardware available but signals need to be set up (confirm) ie. BCMs, etc
  • Z. Ahmed: scaler checks to be done
    • Plug a known pulser in to channels and verify maps
    • Need to verify that the xscaler displays on hcdesk machines, mapping is correct
  • Jure: Event display
    • Two viewers, one for HMS, one for SHMS
    • Jure will check that the viewers work on the hcdesk machines

HMS

  • Remove the Aerogel and place in clean room for repairs [Hamlet]
  • Brad: update wiki with note that Aerogel will not be available
  • Eric Pooser: Jack was involved in pulling the Aerogel -- not sure if Aerogel is out yet

SHMS

Drift Chambers

  • System owners: E. Christy, D. Biswas
  • Now resurveyed
  • No tracks from the analyzer yet, Jure is trying to figure out the PARAM file settings
  • Working on chasing remaining missing signals
  • Plan to do voltage scan and check efficiency with new Eth:Ar mix
  • Need to do drift times and T0 calibration in the software

Hodoscopes

  • System Owners: S. Malace, M. Jones
  • Difficulties with restoring HV this morning (Wed)
  • Simona would like to be the person who manages HVs for hodoscopes for now so she can watch for problems
  • Quartz plane status: stable

Shower/Pre-shower

  • System owners: H. Mkrtchyan, Yerevan
  • Trip currents updated in GUI
  • Please don't change HVs without contacting Hamlet
  • Ask Chuck will turn the Shower fans on -- just leave them on
  • Preshower
    • All channels good
    • Gains are high, can see muons in all channels
    • May need to remove amplifier and replace with passive splitter
      • Be sure to communicate with Brad before doing this
    • Active splitters removed for now
      • Can reinstall passive splitters later (post-KPP)
  • Shower
    • One bad PMT (confirmed)
      • Block is outside acceptance, so maybe fix in summer?

Aerogel Cerenkov

  • System owners: A. Mkrtchyan, T. Horn
  • Splitter was removed
  • All 14 signals are now present (missing channel restored)
    • Good signals observed

NGC

  • System owners: D. Day, E. Pooser
  • Tank will be filled with Argon after it is installed.
  • Waiting for installation into the SHMS stack (presently on the Hall C floor)
  • HV and signal cables are run and verified into DAQ upstairs
  • PMT targets removed 2/16; tank read for installation

HGC

  • System owners: G. Huber, B. Li
  • Will be filled with CO2.
  • Pump and fill is in progress this week (Brad)
  • All channels look alive
  • Hardware and basic software set up is in good shape
  • May want to switch FADCs front-end digitizer FSR to 2V later